# LinPrim: Linear Primitives for Differentiable Volumetric Rendering

### v3

@@ Abstract
Volumetric rendering has become central to modern novel view synthesis methods, which use differentiable rendering to optimize 3D scene representations directly from observed views. While many recent works build on NeRF [18] or 3D Gaussians [13], we explore an alternative volumetric scene representation. More specifically, we introduce two new scene representations based on linear primitives—octahedra and tetrahedra—both of which define homogeneous volumes bounded by triangular faces. To optimize these primitives, we present a differentiable rasterizer that runs efficiently on GPUs, allowing end-to-end gradient-based optimization while maintaining real-time rendering capabilities. Through experiments on real-world datasets, we demonstrate comparable performance to state-of-the-art volumetric methods while requiring fewer primitives to achieve similar reconstruction fidelity. Our findings deepen the understanding of 3D representations by providing insights into the fidelity and performance characteristics of transparent polyhedra and suggest that adopting novel primitives can expand the available design space.  [^1] https://nicolasvonluetzow.github.io/LinPrim/
@@ 1 Introduction
Approaches in NVS utilize a wide range of scene representations that differ significantly from the conventional triangle-mesh-based pipelines used in most 3D applications. While triangle meshes are central to standard 3D content creation, their direct optimization in NVS tasks remains challenging [21, 6]. Instead, it is essential to explore the extensive design space of possible representations to assess how they impact fidelity, performance and stability in NVS settings.
@@
We build on the 3DGS pipeline to accommodate the new primitives. Our method first constructs polyhedra from a concise set of features in a fully differentiable manner. During rendering, we then rely on simple ray-triangle intersections to determine the opacity of each primitive and successively blend them. As a result, we can backpropagate image-space errors through the intersection calculations to distribute gradients onto the geometric features of each polyhedron. The optimization pipeline remains analogous to 3DGS, yet the bounded, triangular geometry of our primitives expands the applicability of differentiable rendering pipelines. Finally, we analyze the photometric reconstruction quality and performance of our approaches on real-world scenes. We are able to show comparable results while retaining real-time rendering speeds.
@@ 3.1 Linear Rendering Primitives
Figure 2: Overview of our method. Primitives are constructed from SfM points. During rendering, primitives are preprocessed according to the camera pose and sorted front-to-back. Afterwards, the list is traversed and the individual color contributions are blended for each pixel. Through comparison with a known view, the primitives’ features and the population are adjusted based on gradient flow.
@@ 3.2 Rendering Process
During preprocessing, we construct each primitive from the aforementioned features (i.e. the position, distances, and rotation). Given the camera pose, we then transform the primitives into the camera’s coordinate system and calculate the view-dependent color based on the viewing direction. Afterwards, we project these primitives into the 3D affine ray space approximation introduced in EWA Splatting [32, 33]. Because small primitives are needed to achieve high-fidelity reconstructions, and because the approximation error grows with increasing distance from the center points, we found that the resulting error remains negligible while still reducing overall performance overhead. Lastly, we compute each primitive’s projected bounding box in screen space to later derive the pixels from which it is visible.
@@
During rasterization, each pixel ray either intersects a given primitive twice or not at all since the introduced primitives are convex. To find intersections, we use the Möller-Trumbore intersection algorithm (MTIA) [19], which is fully differentiable and thus usable in our gradient-based optimization. If there are no intersections, the primitive is not visible; if there are two, the opacity depends on the distance between the intersection points. As our reconstruction is scale ambiguous, we must normalize the obtained intersection distances to ensure consistent appearance across different scene sizes. We do this by determining the smallest distance $d$ of the primitives and then scaling all distance measurements accordingly in a manner similar to EVER [17]. I.e. given the opacity of a primitive $\alpha$, we define the density of an octahedron as:
@@ 3.3 Anti-Aliasing
We adopt their 3D smoothing strategy to limit how small primitives can become based on visibility from the training views. As a result, each primitive is visible from at least one pixel of a training view. However, applying their 2D Mip filter approach is not possible for our primitives because they rely on modifying the screen space Gaussian distribution. Instead, for each primitive, we identify vertices that lie on the bounding box used for tiling and shift them outward, parallel to the screen. This effectively enlarges the bounding box footprint and increases the minimum size at which the primitive appears. Moreover, because the fall-off within each primitive is still treated as linear, expanding the distance from the center raises the opacity along the primitive as it is evaluated over a larger extent. Conceptually, this approach acts as a coarse approximation that serves to band-limit the high frequencies that would otherwise cause visible aliasing. Although this approach is not fully mathematically equivalent to a 2D Mip filter, it captures its function and empirically mirrors its behavior. We evaluate the effectiveness of the filters in Appendix B.
@@ 3.4 Optimization
As a first step, backpropagation through the blending process yields gradients with respect to each primitive’s alpha and color. When these gradients are propagated further, the alpha gradient indicates whether the primitive should be more or less transparent and, consequently, whether the intersections should move closer together or farther apart along the pixel’s ray. By following the resulting intersection gradients, we can adjust each primitive’s center and vertices accordingly to achieve the desired changes. This corresponds to backpropagation through the MTIA, where the intersection depth is influenced by the position of each of the corners of the triangle. In turn, the positions of the corners depend on the position of the ray space center and the offsets to it. Combining the above, by aggregating over the corners and both intersections, we can propagate gradients from alpha to the geometric features.
@@ 3.5 Population Control
Creating new primitives is realized as either cloning or splitting existing ones. For this, primitives that have large view-space positional gradients are chosen, with smaller ones being cloned and larger ones being split based on the size of the primitives, as described above, and the size of the scene. Cloning a primitive creates a new primitive with the same features but without the gradient momentum which is directly applicable to our representations.
Splitting, on the other hand, usually involves creating two new primitives with smaller sizes and placing them by sampling from the original Gaussian as a PDF. To adjust this for octahedra, since the distances are aligned with the coordinate axes and symmetric, we sample the new positions from a PDF with standard deviations set to the corresponding distance and apply rotation afterwards. This way, primitives are more likely to spread along the longer dimensions of the octahedra. For tetrahedra, this is not trivially possible and we instead set all standard deviations to half the largest distance. A sample result of the resulting population of octahedra can be seen in Figure 3.
@@ 4 Experiments
We evaluate our results on Mip-NeRF 360 [2] and ScanNet++ v2 [29]. For Mip-NeRF 360, we use all 9 available scenes, while for ScanNet++, we use only the first five NVS scenes with available test views. We train and test on the official splits provided by each dataset and maintain consistent parameters across all evaluated scenes. All approaches are trained for 30k iterations.
@@
While there is a large body of work offering improvements to 3DGS, we mainly focus on the original work and the changes made by Mip-Splatting, as they most closely relate to the foundational changes made by our approach. We believe that much of the recent progress made on improving 3DGS can similarly be applied to our approach to further improve quality and performance metrics.
@@ 4.1 Reconstruction Quality
We measure reconstruction quality using the standard image metrics PSNR, SSIM and LPIPS. On the ScanNet++ dataset, our method consistently achieves scores close to both 3DGS and Mip-Splatting despite using significantly fewer primitives, as visible in Table 1. We also compare against GS-MCMC [14], which allows precise limitations to the primitive population, and find that our results remain comparable, albeit slightly worse. For Mip-NeRF 360 scenes, we see a similar trend of comparable performance but with a more compact representation, as can be seen in Table 2. These results align with other works that suggest that the number of used Gaussians can be drastically reduced without a negative impact on performance through introduction of more advanced population control paradigms [8, 9]. While our approach introduces fundamental changes, we deliberately preserve components not directly related to the primitives. As such, more advanced paradigms can be applied in our primitives in a similar manner.
Figure 4 illustrates that our method faithfully reconstructs many challenging regions while sometimes showing slightly lower overall fidelity. These outcomes stem, in part, from the bounded nature of our primitives, which enforce sharper, more “binary” decisions in underconstrained areas. In practice, this can lead to crisper reconstructions in reflective or transparent regions—such as glass or distant backgrounds—but can also result in more conspicuous edges on rarely observed surfaces like ceilings. Overall, we find our approach to work the best in smaller, more densely captured scenes, as it retains strong geometric and visual clarity without overly densifying frequently observed regions. Further results on Mip-NeRF 360 scenes can be seen in Appendix C.
Table 1: Quantitative Results on ScanNet++ v2 scenes [29]. We limit GS-MCMC to use as many primitives as our approach does on average. Our method generates results of comparable performance while using much fewer primitives than 3DGS and Mip-Splatting.
PSNR 39f36da05b 5a269ba6fe dc263dfbf0 08bbbdcc3d fb564c935d Prim. 3DGS [13] 27.693 23.675 23.436 22.444 23.210 738k Mip-Splatting [30] 27.728 23.636 23.454 22.570 23.221 977k GS-MCMC [14] 27.902 23.787 23.782 23.657 23.371 255k LinPrim (ours) 27.557 23.845 22.668 22.312 23.801 255k
Table 2: Quantitative Results on Mip-NeRF 360 scenes [2]. Our approach generates results comparable in performance while using fewer primitives.
@@
LinPrim (ours) 26.63 0.803 0.221 1.79M
@@ 4.3 Limitations
While our method demonstrates comparable results and is easily understood and modified, it is not without drawbacks. First, the bounded nature of our primitives can introduce hard edges in poorly observed regions, resulting in more "segment-like" artifacts under limited view coverage. In these regions, the ability of Gaussian kernels to smoothly transition into each other leads to more visually pleasing results, even if the reconstruction is not necessarily better. Second, the current population control processes are not as consistent as their Gaussian counterparts. While we did adapt the processes to our primitives, the underlying paradigms are not as suitable for our polyhedron primitives. More specifically, polyhedra could benefit from more specialized processes for placing newly split primitives and improved criteria for determining which areas should be densified. It would be especially interesting to explore enhancements to 3DGS that address the above issues, as such improvements could disproportionately benefit our polyhedron-based approaches.
@@
Our work provides a new approach to novel view synthesis, which extends the design space of 3D scene representations. It has the potential to improve the quality and performance of 3D reconstruction and downstream applications, benefiting research in graphics, vision, robotics, and related fields. We do not anticipate any adverse environmental, societal or ethical effects.
@@ 5 Conclusion
We have presented novel approaches to volumetric scene reconstruction that leverage linear primitives, specifically octahedra and tetrahedra. Our method relies on a compact set of shape parameters and optimizes both geometry and appearance in a fully differentiable manner. By leveraging the inherent symmetries of octahedra, we achieve stable, high-fidelity results while maintaining a low memory footprint. At the same time, our experiments with tetrahedra demonstrate that the approach generalizes to other triangle-based volumes, underscoring its flexibility. Despite employing fewer primitives, our method consistently achieves comparable performance across challenging datasets, confirming its viability for real-world applications. Additionally, the bounded nature of our primitives yields more pronounced structural decisions in some rarely observed regions, sometimes leading to crisper reconstructions.
Promising future work could focus on leveraging the inherent properties of our primitives. As a concrete example, their triangular faces closely resemeble traditional meshes. Developing efficient and accurate methods to binarize primitive opacity [7] presents significant potential for directly optimizing mesh representations from images. Additionally, although the rendering enhancements introduced by EVER [17] require significant changes for Gaussian-based methods, our primitives inherently allow for easily finding intersections with homogeneous primitives, avoiding the need for ray-tracing and enabling exact blending without additional computational overhead. Finally, integrating more specialized or advanced population control mechanisms could significantly enhance performance by improving primitive density in optimized scenes.

### v4

@@ Abstract
Volumetric rendering has become central to modern novel view synthesis methods, which use differentiable rendering to optimize 3D scene representations directly from observed views. While many recent works build on NeRF [18] or 3D Gaussians [13], we explore an alternative volumetric scene representation. More specifically, we introduce two new scene representations based on linear primitives—octahedra and tetrahedra—both of which define homogeneous volumes bounded by triangular faces. To optimize these primitives, we present a differentiable rasterizer that runs efficiently on GPUs, allowing end-to-end gradient-based optimization while maintaining real-time rendering capabilities. Through experiments on real-world datasets, we demonstrate comparable performance to state-of-the-art volumetric methods while requiring fewer primitives to achieve similar reconstruction fidelity. Our findings deepen the understanding of 3D representations by providing insights into the fidelity and performance characteristics of transparent polyhedra and suggest that adopting novel primitives can expand the available design space.  [^1] Website: https://nicolasvonluetzow.github.io/LinPrim/
@@ 1 Introduction
Approaches in NVS utilize a wide range of scene representations that differ significantly from the conventional triangle-mesh-based pipelines used in most 3D applications. While triangle meshes are central to standard 3D content creation, their direct optimization in NVS tasks remains challenging [21, 6]. Instead, it is essential to explore the extensive design space of possible representations to assess how they impact fidelity, performance, and stability in NVS settings.
@@
We build on the 3DGS pipeline to accommodate the new primitives. Our method first constructs polyhedra from a concise set of features in a fully differentiable manner. During rendering, we then rely on simple ray-triangle intersections to determine the opacity of each primitive and successively blend them. As a result, we can backpropagate image-space errors through the intersection calculations to distribute gradients onto the geometric features of each polyhedron. The optimization remains analogous to 3DGS, yet the bounded, triangular geometry of our primitives expands the applicability of differentiable rendering pipelines. Finally, we analyze the photometric reconstruction quality and performance of our approaches on real-world scenes. We are able to show comparable results while retaining real-time rendering speeds.
@@ 3.1 Linear Rendering Primitives
Figure 2: Overview of our method. Primitives are constructed from SfM points. During rendering, primitives are preprocessed according to the camera pose and sorted front-to-back. Afterward, the list is traversed and the individual color contributions are blended for each pixel. Through comparison with a known view, the primitives’ features and the population are adjusted based on gradient flow.
@@ 3.2 Rendering Process
During preprocessing, we construct each primitive from the aforementioned features (i.e. the position, distances, and rotation). Given the camera pose, we then transform the primitives into the camera’s coordinate system and calculate the view-dependent color based on the viewing direction. Afterward, we project these primitives into the 3D affine ray space approximation introduced in EWA Splatting [32, 33]. Because small primitives are needed to achieve high-fidelity reconstructions, and because the approximation error grows with increasing distance from the center points, we found that the resulting error remains negligible while still reducing overall performance overhead (see Appendix D). Lastly, we compute each primitive’s projected bounding box in screen space to later derive the pixels from which it is visible.
@@
During rasterization, each pixel ray either intersects a given primitive twice or not at all, since the introduced primitives are convex. To find intersections, we use the Möller-Trumbore intersection algorithm (MTIA) [19], which is fully differentiable and thus usable in our gradient-based optimization. If there are no intersections, the primitive is not visible; if there are two, the opacity depends on the distance between the intersection points. As our reconstruction is scale ambiguous, we must normalize the obtained intersection distances to ensure consistent appearance across different scene sizes. We do this by determining the smallest distance $d$ of the primitives and then scaling all distance measurements accordingly in a manner similar to EVER [17]. I.e. given the opacity of a primitive $\alpha$, we define the density of an octahedron as:
@@
We did experiment with optimizing for the unnormalized densities directly instead of the intermediate value of the opacity, but found no performance gains in practice, while decreasing the interpretability of the feature values.
@@ 3.3 Anti-Aliasing
We adopt their 3D smoothing strategy to limit how small primitives can become based on visibility from the training views. As a result, each primitive is visible from at least one pixel of a training view. However, applying their 2D Mip filter approach is not possible for our primitives because they rely on modifying the screen space Gaussian distribution. Instead, for each primitive, we identify vertices that lie on the bounding box used for tiling and shift them outward, parallel to the screen. This effectively enlarges the bounding box footprint and increases the minimum size at which the primitive appears. Moreover, because the fall-off within each primitive is still treated as linear, expanding the distance from the center raises the opacity along the primitive as it is evaluated over a larger extent. Conceptually, this approach acts as a coarse approximation that serves to band-limit the high frequencies that would otherwise cause visible aliasing. Although this approach is not fully mathematically equivalent to a 2D Mip filter, it captures its function and empirically mirrors its behavior. We evaluate the effectiveness of the filters in Appendix C.
@@ 3.4 Optimization
As a first step, backpropagation through the blending process yields gradients with respect to each primitive’s alpha and color. When these gradients are propagated further, the alpha gradient indicates whether the primitive should be more or less transparent and, consequently, whether the intersections should move closer together or farther apart along the pixel’s ray. By following the resulting intersection gradients, we can adjust each primitive’s center and vertices accordingly to achieve the desired changes. This corresponds to backpropagation through the MTIA, where the intersection depth is influenced by the position of each of the corners of the triangle. In turn, the positions of the corners depend on the position of the ray space center and the offsets to it. Combining the above, by aggregating over the corners and both intersections, we can propagate gradients from alpha to the geometric features. Further details are in Appendix E.
@@ 3.5 Population Control
Creating new primitives is realized as either cloning or splitting existing ones. For this, primitives that have large view-space positional gradients are chosen, with smaller ones being cloned and larger ones being split based on the size of the primitives, as described above, and the size of the scene. Cloning a primitive creates a new primitive with the same features but without the gradient momentum, which is directly applicable to our representations.
Splitting, on the other hand, usually involves creating two new primitives with smaller sizes and placing them by sampling from the original Gaussian as a PDF. To adjust this for octahedra, since the distances are aligned with the coordinate axes and symmetric, we sample the new positions from a PDF with standard deviations set to the corresponding distance and apply rotation afterwards. This way, primitives are more likely to spread along the longer dimensions of the octahedra. For tetrahedra, this is not trivially possible, and we instead set all standard deviations to half the largest distance. A sample result of the resulting population of octahedra can be seen in Figure 3.
LinPrim-MCMC.
We aim to keep as many factors - apart from the primitives themselves - fixed between our approach and the Gaussian baseline. As such, our population control paradigms closely mirror the behavior of 3DGS. Nonetheless, we demonstrate the fact that LinPrim remains compatible with advancements made on Gaussian kernels and analyze the impact of population control by implementing a separate population control technique from GS-MCMC [14] for our octahedron primitives. They rethink population control as drawing samples from the distribution of the scene, which not only improves performance but also allows precise control over the final population size.
We again adopt their approach as close as possible to minimize undesired performance impacts. In this case, we essentially act as if our primitives are Gaussians with the same features. We account for the differences in magnitude between octahedron distances and Gaussian standard deviations by scaling down distances by a factor of $2.6$. After this scaling, the size of the supposed Gaussians more closely resembles the primitives they imitate. A related concept was explored as Distribution Alignment in Beyond Gaussians [4].
@@ 4 Experiments
We evaluate our results on Mip-NeRF 360 [2] and ScanNet++ v2 [29]. For Mip-NeRF 360, we use all 9 available scenes, while for ScanNet++, we use only the first five NVS scenes with available test views. Mip-NeRF 360 primarily features forward-facing captures centered around a single object or region of interest, whereas ScanNet++ consists of indoor scenes with more diverse camera trajectories and test views located farther from the training coverage. We train and test on the official splits provided by each dataset and maintain consistent parameters across all evaluated scenes. All approaches are trained for 30k iterations.
@@
While a large body of work has proposed improvements to 3DGS, we primarily focus on the original formulation, the modifications introduced by Mip-Splatting, and the convex primitives from 3D Convex Splatting (3DCS) [11], as these are most closely aligned with the foundational changes introduced in our method. It is important to note that, although most evaluated approaches share an equal number of parameters per primitive, each 3D convex in 3DCS is defined by six points, resulting in a higher overall parameter count and making a direct comparison of parameter budgets less straightforward.
@@ 4.1 Reconstruction Quality
We measure reconstruction quality using the standard image metrics PSNR, SSIM, and LPIPS. On the ScanNet++ v2 dataset, our method consistently achieves scores close to 3DGS, Mip-Splatting, and 3DCS despite using significantly fewer primitives, as visible in Table 1. We also compare against GS-MCMC [14], which enables explicit control over the primitive population, and demonstrate that our primitives can effectively leverage the same population heuristics, achieving improved average performance at equal population counts. For Mip-NeRF 360 scenes, we see a similar trend of comparable performance but with a more compact representation, as can be seen in Table 2. These results align with other works that suggest that the number of used Gaussians can be drastically reduced without a negative impact on performance through the introduction of more advanced population control paradigms [8, 9].
Figure 4 illustrates that our method faithfully reconstructs many challenging regions while sometimes showing slightly lower overall fidelity. These outcomes stem, in part, from the bounded nature of our primitives, which enforce sharper, more “binary” decisions in underconstrained areas. In practice, this can lead to better depth estimates and crisper reconstructions in reflective or transparent regions—such as glass or distant backgrounds—but can also result in more conspicuous edges on rarely observed surfaces like ceilings (see Appendices F and F). Overall, we find our approach to work the best in smaller, more densely captured scenes, as it retains strong geometric and visual clarity without overly densifying frequently observed regions. Further results on Mip-NeRF 360 scenes can be seen in Appendix H.
Table 1: Quantitative Results on ScanNet++ v2 scenes [29]. We limit the MCMC-based approaches to use as many primitives as LinPrim and 3DGS do on average. Our method achieves comparable performance while using much fewer primitives under default population control, and slightly higher quality when combined with MCMC-based densification. Scene identifiers are shortened for readability.
39f36d 5a269b dc263d 08bbbd fb564c Mean Prim. 3DGS [13] 27.70 23.68 23.44 22.44 23.21 24.09 738k Mip-Splatting [30] 27.73 23.64 23.45 22.57 23.22 24.12 977k 3DCS [11] 27.59 24.19 22.92 22.78 23.84 24.26 440k LinPrim 27.56 23.84 22.67 22.31 23.80 24.04 255k GS-MCMC [14] 27.90 23.79 23.78 23.66 23.37 24.50 255k 27.70 23.40 23.80 23.44 22.28 24.12 738k LinPrim + MCMC 27.62 24.03 23.33 22.98 24.07 24.41 255k 27.74 24.13 23.57 23.33 23.97 24.55 738k
Table 2: Quantitative Results on Mip-NeRF 360 scenes [2]. We limit the MCMC-based approaches to use as many primitives as 3DGS do on average. † 3DCS [11] uses two different sets of hyperparameters for outdoor and indoor scenes.
@@
3DCS [11] 27.22 0.801 0.208 1.02M
LinPrim 26.63 0.803 0.221 1.79M
GS-MCMC [14] 28.09 0.836 0.187 3.32M
LinPrim + MCMC 27.04 0.812 0.211 3.32M
@@ 4.3 Limitations
While our method demonstrates comparable results and is easily understood and modified, it is not without drawbacks. First, the bounded nature of our primitives can introduce hard edges in poorly observed regions, resulting in more "segment-like" artifacts under limited view coverage. In these regions, the ability of Gaussian kernels to smoothly transition into each other leads to more visually pleasing results, even if the reconstruction is not necessarily better. Second, while our adapted population control and MCMC-based updates modulate primitive populations well, these processes were originally designed for Gaussian primitives and are therefore not yet fully optimal for polyhedral representations. Future work could explore more specialized strategies for splitting, cloning, and sampling that better exploit the geometric structure of polyhedra. Finally, the rendering efficiency of our current implementation remains below that of the optimized Gaussian-based rasterizer (see Appendix A). Reducing the computational overhead of polygon-based rasterization by leveraging existing optimized tools and pipelines for triangle-based rendering would substantially enhance the scalability and practicality of our approach.
@@
Our work provides a new approach to novel view synthesis, which extends the design space of 3D scene representations. It has the potential to improve the quality and performance of 3D reconstruction and downstream applications, benefiting research in graphics, vision, robotics, and related fields. We do not anticipate any adverse environmental, societal, or ethical effects.
@@ 5 Conclusion
We introduced a novel framework for volumetric scene reconstruction based on linear primitives, specifically octahedra and tetrahedra. Our formulation jointly optimizes geometry and appearance through a compact set of shape parameters in a fully differentiable manner. Exploiting the inherent symmetries of octahedra enables stable, high-fidelity reconstructions while maintaining a low memory footprint. In addition, our experiments with tetrahedra show that the concept generalizes naturally to other triangle-based volumes, highlighting its flexibility. Despite using significantly fewer primitives, our method achieves competitive performance across challenging datasets, demonstrating its effectiveness for real-world applications. At the same time, the bounded nature of our primitives promotes more distinct structural representations, resulting in crisper reconstructions in highly view-dependent areas and a more accurate geometric understanding, all without increasing sensitivity to initialization (see Appendices B and F).
Promising future work could focus on further leveraging the inherent properties of our primitives. For instance, their triangular faces closely resemble traditional meshes, suggesting a natural bridge between primitive-based and mesh-based representations. Developing efficient and accurate methods to binarize primitive opacity [7] could enable direct mesh optimization from images. Moreover, while the rendering enhancements introduced by EVER [17] require substantial modifications for Gaussian-based methods, our primitives naturally support intersection tests with homogeneous volumes, eliminating the need for explicit ray tracing and allowing exact blending without additional computational overhead. Leveraging existing triangle-mesh rendering pipelines could further accelerate our approach, substantially improving efficiency while removing the need for ray-space approximations (see Appendices A and D). Beyond static reconstruction, exploring downstream applications of our primitives—such as dynamic or deformable scene representations—constitutes an interesting direction for future research.